# Asymptotics of the maximum likelihood estimator of the location parameter of Pearson Type VII distribution

### v1

@@ 1. Introduction
Thus we obtain (7.2) and (7.3) and the proof is completed.
@@
The variance of the maximum likelihood estimator of the parameter $m$ was dealt with by Taylor’s unpublished manuscript [14].
8. Numerical computations
@@
We perform simulation studies using the software R to illustrate the properties of the maximum likelihood estimator. We used R version 4.5.1. We deal with $nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]$ appearing in Theorem 1.6. We can assume that $\theta=0$ without loss of generality. We consider parameters $m=0.1\times k$ for $6\leq k\leq 15$ and sample sizes $n=10,50,100,500,1000$. In each choice of the pair $(m,n)$, we compute the average of $n\left(\hat{\theta}_{n}\right)^{2}$ over $10^{7}$ samples of size $n$ generated by the rpearsonVII() function in the package ‘PearsonDS’, and in the optimization, we use the nlminb() function with the starting point being the median. Table 1 gives the result.
@@
In the case of $m=1$, [6, Table 2] gives numerical computations for $n=5,6,\dots,14,15,20,25,\dots,35,40$. This is consistent with the numerical computation in [1, Table 2] for the joint estimation of the location and scale.
The convergence becomes faster as the parameter $m$ increases. Since $\hat{\theta}_{n}$ appears not to be square-integrable if $m=0.6$ and $n=10$, and the numerical computation is not stable although $\hat{\theta}_{n}$ appears to be square-integrable if $m=0.7$ and $n=10$, we mark them as not available (NA). By this table, we conjecture that for each $m$, $\left(nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]\right)_{n}$ is decreasing in $n$.
[…]
0.6 NA 25.756 16.108 13.756 13.545 13.333
0.7 NA 7.236 6.566 6.160 6.116 6.071
0.8 9.046 4.156 3.935 3.783 3.766 3.750
0.9 4.524 2.832 2.729 2.655 2.648 2.639
1 2.908 2.108 2.052 2.011 2.006 2.000
1.1 2.103 1.659 1.624 1.599 1.595 1.591
1.2 1.630 1.356 1.332 1.314 1.311 1.310
1.3 1.320 1.138 1.123 1.108 1.107 1.106
1.4 1.105 0.977 0.964 0.954 0.953 0.952
1.5 0.945 0.851 0.842 0.835 0.834 0.833
Table 1. Simulated values of $n\,\mathbb{E}\!\left[(\hat{\theta}_{n}-\theta)^{2}\right]$ ($\theta=0$). Rows correspond to $m$; columns to the sample size $n$. The column labeled $\infty$ reports the theoretical limit (m+1)/(m(2m1))(m+1)/(m(2m-1)) given by Theorem 1.6. We round the results to three decimal places.

### v2

@@ 1. Introduction
Thus we obtain (7.2) and (7.3), and the proof of Theorem 1.6 is completed. We provide numerical evidence for Theorem 1.6 in Subsection 8.4 below.
@@
(i) The variance of the maximum likelihood estimator of the parameter $m$ was dealt with by Taylor’s unpublished manuscript [20]. [14, pp.396–399] discusses the parameter estimation other than the location.
(ii) By symmetry, we strongly expect that $\hat{\theta}_{n}$ is unbiased, that is, $E\left[\hat{\theta}_{n}\right]=\theta$ for each $n$ such that $\hat{\theta}_{n}$ is integrable. However, to the best of our knowledge, no rigorous proof of this fact has been established. We provide numerical evidence for this in Subsection 8.1 below.
8. Simulation study
@@
We perform simulation studies using R to illustrate the properties of the maximum likelihood estimator. We consider parameter values $m=0.1\cdot k$ for $6\leq k\leq 15$ and sample sizes $n=10,50,100,500,1000$. In the tables below, the rows correspond to $m$ and the columns to the sample size $n$, unless otherwise stated. In the simulations, we let $\theta=0$. We use R version 4.5.2. For each pair $(m,n)$, we generate $N=10^{7}$ samples of size $n$ using the rpearsonVII() function in the package ‘PearsonDS’, and compute the corresponding MLEs. For the optimization, we use the nlminb() function with the sample median as the starting value.
8.1. Bias
We compute $\left|E\left[\hat{\theta}_{n}-\theta\right]\right|$. We approximate $E\left[\hat{\theta}_{n}-\theta\right]$ by $\displaystyle\frac{1}{N}\sum_{i=1}^{N}Z_{i}$, where $(Z_{i})_{i}$ are i.i.d. random variables distributed as $\hat{\theta}_{n}-\theta$.
$\n\displaystyle m\backslash$ n 10 50 100 500 1000
0.6 2.3 $3.7\cdot 10^{-5}$ $8.5\cdot 10^{-5}$ $1.1\cdot 10^{-4}$ $4.4\cdot 10^{-5}$
0.7 $7.8\cdot 10^{-4}$ $1.7\cdot 10^{-5}$ $1.2\cdot 10^{-4}$ $2.6\cdot 10^{-5}$ $2.9\cdot 10^{-6}$
0.8 $3.3\cdot 10^{-5}$ $2.9\cdot 10^{-5}$ $1.9\cdot 10^{-5}$ $3.2\cdot 10^{-5}$ $5.6\cdot 10^{-6}$
0.9 $7.3\cdot 10^{-5}$ $7.1\cdot 10^{-5}$ $5.2\cdot 10^{-5}$ $7.2\cdot 10^{-6}$ $3.9\cdot 10^{-5}$
1.0 $1.8\cdot 10^{-4}$ $5.2\cdot 10^{-5}$ $6.2\cdot 10^{-5}$ $4.5\cdot 10^{-6}$ $2.3\cdot 10^{-5}$
1.1 $8.6\cdot 10^{-6}$ $6.5\cdot 10^{-5}$ $2.0\cdot 10^{-5}$ $5.2\cdot 10^{-6}$ $1.2\cdot 10^{-5}$
1.2 $6.0\cdot 10^{-5}$ $6.9\cdot 10^{-5}$ $4.3\cdot 10^{-5}$ $1.4\cdot 10^{-5}$ $7.6\cdot 10^{-6}$
1.3 $3.7\cdot 10^{-6}$ $3.0\cdot 10^{-6}$ $8.9\cdot 10^{-5}$ $4.5\cdot 10^{-6}$ $3.2\cdot 10^{-6}$
1.4 $5.7\cdot 10^{-5}$ $1.8\cdot 10^{-5}$ $2.7\cdot 10^{-5}$ $6.3\cdot 10^{-6}$ $4.7\cdot 10^{-6}$
1.5 $4.2\cdot 10^{-5}$ $2.8\cdot 10^{-5}$ $6.3\cdot 10^{-6}$ $1.6\cdot 10^{-5}$ $1.5\cdot 10^{-6}$
Table 1. Simulated values of $\left|E\left[\hat{\theta}_{n}-\theta\right]\right|$.
This table suggests that $\hat{\theta}_{n}$ is not integrable for $m=0.6$ and $n=10$. In this case, it is interesting to compute $P\left(\hat{\theta}_{n}-\theta>r\right)$ for large $r$. We approximate this probability by the proportion of simulated values which are larger than $r$ among the $N=10^{7}$ simulated observations.
$r$ $10^{3}$ $5\cdot 10^{3}$ $10^{4}$ $5\cdot 10^{4}$ $10^{5}$ $5\cdot 10^{5}$ $10^{6}$
$-1.22$ $-1.08$ $-1.02$ $-0.94$ $-0.94$ $-0.95$ $-0.96$
Table 2. Simulated values of $\dfrac{\log P(\hat{\theta}_{n}-\theta>r)}{\log r}$.
This table suggests that there exists $C>0$ such that $P\left(\hat{\theta}_{n}-\theta>r\right)\geq Cr^{-1}$ for large $r$, which implies $\hat{\theta}_{n}$ is not integrable.
8.2. Asymptotic normality
Recall that $\varphi_{m}$ is the density of the normal distribution $N\left(0,\dfrac{m+1}{m(2m-1)}\right)$. We consider the Kolmogorov–Smirnov metric:
$\Delta_{n}\coloneqq\sup_{x\in\mathbb{R}}\left|P\left(\hat{\theta}_{n}-\theta\leq x\right)-\int_{-\infty}^{x}\varphi_{m}(t)dt\right|.$
We approximate $P\left(\hat{\theta}_{n}-\theta\leq x\right)$ by the empirical cumulative distribution function $\displaystyle\frac{1}{N}\sum_{i=1}^{N}{\bf 1}_{\{Z_{i}\leq x\}}$ where $(Z_{i})_{i}$ are i.i.d. random variables distributed as $\hat{\theta}_{n}-\theta$. We recall the Dvoretzky-Kiefer-Wolfowitz-Massart inequality [9, Theorem 1.32]:
$P\left(\sup_{x\in\mathbb{R}}\left|\frac{1}{N}\sum_{i=1}^{N}{\bf 1}_{\{Z_{i}\leq x\}}-P\left(\hat{\theta}_{n}-\theta\leq x\right)\right|>\epsilon\right)\leq 2\exp(-2N\epsilon^{2}).$
$\n\displaystyle m\backslash$ n 10 50 100 500 1000
0.6 $1.1\cdot 10^{-1}$ $2.4\cdot 10^{-2}$ $1.2\cdot 10^{-2}$ $2.4\cdot 10^{-3}$ $1.3\cdot 10^{-3}$
0.7 $5.3\cdot 10^{-2}$ $1.1\cdot 10^{-2}$ $5.5\cdot 10^{-3}$ $1.2\cdot 10^{-3}$ $5.8\cdot 10^{-4}$
0.8 $3.4\cdot 10^{-2}$ $6.9\cdot 10^{-3}$ $3.5\cdot 10^{-3}$ $7.9\cdot 10^{-4}$ $4.7\cdot 10^{-4}$
0.9 $2.5\cdot 10^{-2}$ $5.0\cdot 10^{-3}$ $2.5\cdot 10^{-3}$ $6.4\cdot 10^{-4}$ $3.5\cdot 10^{-4}$
1.0 $1.9\cdot 10^{-2}$ $4.0\cdot 10^{-3}$ $2.0\cdot 10^{-3}$ $6.4\cdot 10^{-4}$ $2.9\cdot 10^{-4}$
1.1 $1.5\cdot 10^{-2}$ $3.1\cdot 10^{-3}$ $1.5\cdot 10^{-3}$ $4.2\cdot 10^{-4}$ $3.1\cdot 10^{-4}$
1.2 $1.3\cdot 10^{-2}$ $2.6\cdot 10^{-3}$ $1.4\cdot 10^{-3}$ $3.0\cdot 10^{-4}$ $2.7\cdot 10^{-4}$
1.3 $1.1\cdot 10^{-2}$ $2.3\cdot 10^{-3}$ $1.2\cdot 10^{-3}$ $3.4\cdot 10^{-4}$ $3.2\cdot 10^{-4}$
1.4 $9.3\cdot 10^{-3}$ $1.9\cdot 10^{-3}$ $9.7\cdot 10^{-4}$ $2.8\cdot 10^{-4}$ $2.1\cdot 10^{-4}$
1.5 $8.2\cdot 10^{-3}$ $1.6\cdot 10^{-3}$ $9.3\cdot 10^{-4}$ $3.8\cdot 10^{-4}$ $2.4\cdot 10^{-4}$
Table 3. Simulated values of $\Delta_{n}$.
8.3. Confidence intervals
We consider the pivotal quantity $\sqrt{n}\left(\hat{\theta}_{n}-\theta\right)$ of the model. Let $z_{\beta}$ denote the upper $\beta$-quantile, that is, the value satisfying $P\left(\sqrt{n}(\hat{\theta}_{n}-\theta)\geq z_{\beta}\right)=\beta$ for $0<\beta<\dfrac{1}{2}$. We report the values of $z_{\alpha/2}$ for $\alpha=0.1,0.05,0.01$. We approximate $z_{\alpha/2}$ by sorting $N=10^{7}$ MLEs and using the quantile function in R. Using $z_{\alpha/2}$, we obtain the $100(1-\alpha)\%$ confidence interval for the location parameter $\theta$:
$\left[\hat{\theta}_{n}-\frac{z_{\alpha/2}}{\sqrt{n}},\hat{\theta}_{n}+\frac{z_{\alpha/2}}{\sqrt{n}}\right].$
In the following tables, the column labeled $\infty$ reports the upper $\alpha/2$ quantile of the limiting normal distribution $N\left(0,\dfrac{m+1}{m(2m-1)}\right)$ given by Theorem 1.2.
$\n\displaystyle m\backslash$ n 10 50 100 500 1000 $\infty$
0.6 19.60 7.06 6.46 6.08 6.05 6.01
0.7 6.34 4.34 4.19 4.08 4.06 4.05
0.8 4.09 3.33 3.25 3.20 3.19 3.19
0.9 3.17 2.75 2.71 2.68 2.68 2.67
1.0 2.64 2.38 2.35 2.33 2.33 2.33
1.1 2.29 2.11 2.10 2.08 2.08 2.07
1.2 2.04 1.91 1.90 1.89 1.88 1.88
1.3 1.85 1.75 1.74 1.73 1.73 1.73
1.4 1.70 1.62 1.61 1.61 1.61 1.61
1.5 1.58 1.52 1.51 1.50 1.50 1.50
Table 4. Simulated values of $z_{\alpha/2}$ ($\alpha=0.1$).
$\n\displaystyle m\backslash$ n 10 50 100 500 1000 $\infty$
0.6 37.96 9.14 7.96 7.29 7.22 7.16
0.7 9.46 5.34 5.07 4.87 4.85 4.83
0.8 5.57 4.04 3.91 3.82 3.80 3.80
0.9 4.12 3.33 3.25 3.20 3.19 3.18
1.0 3.35 2.87 2.82 2.78 2.78 2.77
1.1 2.87 2.54 2.51 2.48 2.48 2.47
1.2 2.53 2.29 2.27 2.25 2.25 2.24
1.3 2.28 2.10 2.08 2.06 2.06 2.06
1.4 2.09 1.94 1.93 1.92 1.92 1.91
1.5 1.93 1.82 1.80 1.79 1.79 1.79
Table 5. Simulated values of $z_{\alpha/2}$ ($\alpha=0.05$).
$\n\displaystyle m\backslash$ n 10 50 100 500 1000 $\infty$
0.6 162.34 15.48 11.47 9.73 9.57 9.41
0.7 21.55 7.66 6.91 6.45 6.39 6.35
0.8 10.48 5.58 5.27 5.04 5.01 4.99
0.9 6.99 4.53 4.35 4.21 4.20 4.18
1.0 5.31 3.87 3.75 3.67 3.65 3.64
1.1 4.35 3.41 3.33 3.27 3.26 3.25
1.2 3.74 3.07 3.01 2.96 2.95 2.95
1.3 3.29 2.80 2.75 2.72 2.71 2.71
1.4 2.97 2.59 2.55 2.52 2.52 2.51
1.5 2.71 2.41 2.38 2.36 2.35 2.35
Table 6. Simulated values of $z_{\alpha/2}$ ($\alpha=0.01$).
In the case $m=1$, [11, Section 4] provides a table of quantiles $z_{\alpha/2}$ for sample sizes $n=5,10,\dots,40$ and $\alpha=0.1,0.05,0.01$.
8.4. Asymptotic efficiency
We consider the quantity $nE\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]$ appearing in Theorem 1.6. The column labeled $\infty$ reports the theoretical limit $\dfrac{m+1}{m(2m-1)}$ given by Theorem 1.6.
$\n\displaystyle m\backslash$ n 10 50 100 500 1000 $\infty$
0.6 NA 25.756 16.108 13.756 13.537 13.333
0.7 NA 7.231 6.561 6.162 6.114 6.071
0.8 8.979 4.154 3.933 3.786 3.769 3.750
0.9 4.511 2.831 2.730 2.657 2.648 2.639
1.0 2.907 2.110 2.053 2.011 2.005 2.000
1.1 2.106 1.660 1.623 1.598 1.595 1.591
1.2 1.631 1.356 1.332 1.314 1.311 1.310
1.3 1.322 1.138 1.122 1.108 1.107 1.106
1.4 1.105 0.976 0.964 0.955 0.953 0.952
1.5 0.946 0.851 0.842 0.835 0.834 0.833
Table 7. Simulated values of $n\,\mathbb{E}\!\left[\left(\hat{\theta}_{n}-\theta\right)^{2}\right]$.
@@ 8.4. Asymptotic efficiency
In the case $m=1$, [11, Table 2] provides numerical values for $n=5,6,\dots,14,15,20,25,\dots,35,40$. These results are consistent with the numerical results in [2, Table 2] for the joint estimation of the location and scale.
Acknowledgements. The author would like to express his gratitude to the referees for their helpful comments and suggestions.